Three types of ticketts for a concert, type A at £75, type B at £55 and type C at £30. There are three times as many of type B sold compared to type A. In total 23000 tickets are sold making revenue of £870,000. How many of each type ticket are sold.
Im presuming A=3B in any equation.

Answer by ikleyn(52780)

.

Let x be the number of the A-type tickets.


Then the number of the B-type tickets is 3x, as stated.


Finally, the number of the C-type tickets is  (23000 - x - 3x) = (23000-4x).


Then the money equation (=the revenue equation) is


    75x + 55*(3x) + 30*(23000-4x) = 870000.


Thus I reduced the original problem to one single equation in one unknown.


I leave it to you to solve it.


When you find x from the equation, you will be in position to answer the problem's question.

------------

At this point, I completed my instructions.

The rest is on you.


```````````

The major lesson for you from my post is to learn  HOW  I reduced the original problem to a single equation in one unknown.


At this point,  I open to accept your  "DEEPEST  THANKS"  for my teaching,  and  I  expect to get it  SOON.